OpenStudy (anonymous):

f(x)=2sec^2(x^7). Find f'(x).

OpenStudy (anonymous):

\[=2\frac{ d }{ dx }(secx^7)^2\] \[=2*2(secx^7)*(secx^7)'\]

OpenStudy (anonymous):

\[=4secx^7*\sec(x^7)\tan(x^7)*(x^7)'\]

OpenStudy (anonymous):

\[=4secx^7secx^7tanx^7*7x^6\]

OpenStudy (anonymous):

what did i do wronggg?